Given 5x+2y-6=0 write equation of line parallel to it going through point (-2,-5)
Answer by fractalier(6550)

Find the slope of this line first.
5x + 2y - 6 = 0
2y = -5x + 6
y = (-5/2)x + 3
So you need a slope of -5/2.
Now use the point given and y = mx + b to find b...
-5 = (-5/2)(-2) + b
-5 = 5 + b
b = -10
so your line is
y = (-5/2)x - 10